STUDY OF STATE DENSITY EFFECTS BY IMPULSIVE COLLISIONS OF HIGHLY VIBRATIONALLY EXCITED MOLECULES AND H_{2}O

\underline{Q. LIU}, D. K. HAVEY and A. S. MULLIN, Department of Chemistry and Biochemistry, University of Maryland, College Park, MD.

We investigated how donor molecules’ state density affect impulsive collisions by studing collisional energy transfer of vibrationally excited pyridine-d_{5} (donor) and H_{2}O (bath), and the result was compared with experiments of pyridine:H_{2}O, picoline:H_{2}O and lutidine:H_{2}O. Vibrationally excited pyridine-d_{5} was prepared by absorbing pulsed 266nm UV light from forth harmonic generator of a Nd:YAG laser. A tunable F-center laser with 0.0003 cm^{−1} resolution was used to probe individual H_{2}O (000 to 001) rotational states. The population of each H_{2}O rotational state was measured by time dependent transient IR absorption and the rate of single collision between pyridine-d_{5} and H_{2}O was determined afterwards. The translational energy gain in H_{2}O was obtained by Doppler-broadened linewidths for each rotational state and the rotational energy gain was measured by a population distribution of H_{2}O rotational states with E_{rot} between 1000 and 2000 cm^{−1}. The probability distributions for the large energy transfer of pyridine-d_{5} and H_{2}O was calculated and the result was compared with other donors and H_{2}O.